### v1

# Learning Koopman Dynamics for Safe Legged Locomotion with Reinforcement Learning-based Controller

Abstract
Learning-based algorithms have demonstrated impressive performance in agile locomotion of legged robots. However, learned policies are often complex and opaque due to the black-box nature of learning algorithms, which hinders predictability and precludes guarantees on performance or safety. In this work, we develop a novel safe navigation framework that combines Koopman operators and model-predictive control (MPC) frameworks. Our method adopts Koopman operator theory to learn the linear evolution of dynamics of the underlying locomotion policy, which can be effectively learned with Dynamic Mode Decomposition (DMD). Given that our learned model is linear, we can readily leverage the standard MPC algorithm. Our framework is easy to implement with less prior knowledge because it does not require access to the underlying dynamical systems or control-theoretic techniques. We demonstrate that the learned linear dynamics can better predict the trajectories of legged robots than baselines. In addition, we showcase that the proposed navigation framework can achieve better safety with less collisions in challenging and dense environments with narrow passages.
I INTRODUCTION
Recent advances in reinforcement learning have led to significant improvements in robust and agile quadrupedal locomotion [1, 2, 3, 4, 5, 6]. Extensive experiments have demonstrated robust performance when traversing complex terrains [2, 3, 5], with the potential for efficient training fueled by large computational resources [1] or motion priors [4, 5, 6]. However, learned locomotion policies are often limited to tracking the given velocity command, leaving navigation decisions to a high-level policy.
Some learning-based frameworks also tackle navigation by either training a dedicated high-level policy [7, 8, 9] or developing an end-to-end policy using diverse navigation maps and additional navigation rewards [10, 11, 12]. While these approaches are effective, they demand significant computational resources and extensive user expertise for trail-and-error-based parameter tuning. Moreover, learning-based approaches often demonstrate degraded performance in unseen environments. As such, these approaches lack strong safety guarantees despite appearing to be safe under specific experimental conditions.
Unlike learning-based approaches, conventional model-based approaches tend to integrate obstacle avoidance constraints with model predictive control (MPC) for safe robot navigation. However, these approaches assume explicit knowledge of the robot’s underlying dynamical model. This model tends to be highly complex due to the hybrid dynamics of frequent making and breaking of contacts [13], and highly nonlinear whole-body dynamics [14]. To circumvent such complexity and facilitate constrained optimization, existing methods resort to model simplifications [13, 15, 16], but sacrifice accuracy in the process.
We contribute an easy-to-use framework (Fig. 1 for an overview) for safe navigation that can be used in conjunction with any locomotion controller, learned or otherwise. Our key idea is to approximately encode the closed-loop dynamics of legged motion using a unified linear model that captures the co-evolution of state variables (e.g., pose and velocity) over time. In particular, we use Dynamic Mode Decomposition (DMD) [17] to efficiently learn unified linear models from data. We also propose the use of approximated Koopman operator [18] in conjunction with DMD to encode underlying nonlinear dynamics using a linear system defined in a higher-dimensional space.
Exploiting the fact our learned dynamics are linear, we employ a conventional model-predictive controller (MPC) to generate the desired velocity commands for collision-free navigation in dense and narrow environments. Because the dynamic constraints are linear, our framework is simple and can be used with modest computational resources. We do not assume any knowledge of the underlying dynamics or its structure, and our framework can learn from offline locomotion data (e.g., rollouts of pre-trained controllers with random control commands).
Figure 1: Our approach learns to approximately represent the high-level closed-loop dynamics of a given learning-based locomotion controller using a unified linear model, which we then utilize to ensure model-based safety.
We conducted extensive experiments to evaluate the effectiveness of our framework. First, we compared the long-term prediction accuracy of our unified linear and koopman-based models against that of a component-based linear model [19], and a simple linear integrator model. Our results indicate that both our models consistently outperform the baselines in terms of long-term prediction accuracy on the withheld dataset, with our Koopman-based model performing better than our unified linear model.
Second, we compared each model’s ability to enable safe navigation when used in conjunction with MPC (See Fig. 2 for examples). Specifically, we conducted our evaluation on three different maps of varying difficulty. Our results suggest that our models can enable safe navigation across all the maps, outperforming the component-wise linear and the simple integrator models.
The paper is organized as follows: In Sec. II, we discuss related works and contextualize our contributions. In Sec. III, we introduce DMD and Koopman Operator theory. In Sec. IV, we detail our framework, including offline data collection, model learning, and model-predictive safety. In Sec. V, we discuss experimental results. In Sec. VI, we summarize our contributions and suggest future directions.
(a) Narrow Corridor
(b) Maze
Figure 2: Our approach enables safe navigation in challenging environments with dense obstacles and narrow passages.
II Related Works
II-A Safe Legged Locomotion
It is important to ensure safety during legged robots’ dynamic motions with underactuated bases. Previously, safety has been widely discussed within the context of model-based control [20, 14]. Recent advances in learning-based locomotion [21, 1] frameworks have demonstrated impressive robustness, but there are only limited works that discuss the interpretability [22] and guarantee of safety [23]. A common approach in this context is to decouple a low-level locomotion controller and a high-level navigation policy, where a high-level policy adopts conventional methods such as model predictive control and control barrier functions [24, 25, 26]. Li et al. [19] propose a method that leverages low-dimensional linear models identified from bipedal locomotion policies trained via reinforcement learning. Liao et al. [27] models robots and obstacles as polytopes [28] to enable narrow space navigation. However, they require access to a low-level whole-body controller, which is not available in many real-world scenarios. Instead of using analytic formulations, there exist works to model low-level behaviors using neural networks [29, 30]. However, the learned dynamics are often highly nonlinear, which makes it infeasible to adopt conventional optimization methods and requires additional techniques, such as an informed trajectory sampler [30].
II-B Koopman Operator for Robotics
The Koopman operator theory [31], which offers a linear perspective on nonlinear dynamical systems, has gained growing interest in robotics for its potential to simplify complex dynamical systems [32] and high sample efficiency [33]. They have been successfully utilized in various domains such as spherical robots [32], drones [34, 35, 36], and dexterous manipulation [33, 37, 38]. Leveraging their linearity, Koopman operator methods have widely been combined with model predictive controls to obtain desired command tracking [32, 39], enforce better stability, or obtain safety. The Koopman operator has also proven useful in learning the reference dynamics from trajectory data, which demonstrates superior sample efficiency compared to deep neural networks in dexterous manipulation [33]. In this work, we learn the high-level dynamics of the velocity-tracking policy and leverage a model predictive control framework to obtain safe commands to navigate within environments with obstacles.
III Preliminaries: DMD and Koopman Operators
We begin by providing a brief introduction to DMD and Koopman operator theory [18].
Koopman Representation: Consider a discrete-time autonomous nonlinear dynamical system $\mathrm{x}_{t+1}=F(\mathrm{x}_{t}),$ where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state at time $t$, and $F(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a nonlinear transition function. To precisely represent this nonlinear dynamical system as a linear system, we introduce a set of observables using the so-called lifting function $\psi:\mathcal{X}\rightarrow\mathcal{O}$, where $\mathcal{O}$ is the space of observables. We can now define the Koopman Operator $\mathcal{K}$, an infinite-dimensional operator on the lifting function $\psi(\cdot)$ for the discrete-time nonlinear system as follows
$[\mathcal{K}\psi]=\psi(F(\mathrm{x}_{t}))=\psi(\mathrm{x}_{t+1})$ (1)
If the observables belong to a vector space, $\mathcal{K}$ can be seen as an infinite-dimensional linear map that describes the evolution of the observables: $\psi(\mathrm{x}_{t+1})=\mathcal{K}\psi(\mathrm{x}_{t})$.
In practice, we do not benefit from this representation since it is infinite-dimensional. However, we can approximate $\mathcal{K}$ using a matrix $K\in\mathbb{R}^{p\times p}$ and define a finite set of observables $\psi(t)\in\mathbb{R}^{p}$. Thus, we can rewrite the relationship as $\psi(\mathrm{x}_{t+1})=K\psi(\mathrm{x}_{t})+r(\mathrm{x}_{t}),$ where $r(\mathrm{x}_{t})\in\mathbb{R}^{p}$ is the residual error caused by the finite dimensional approximation, which can be arbitrarily reduced based on the choice of the lifting function, $p$ and $\psi(\cdot)$.
Further, let’s consider a controlled dynamical system $\mathrm{x}_{t+1}=F(\mathrm{x}_{t},\mathrm{u}_{t}),$ where $\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$. Now we aim to satisfy $\psi(\mathrm{x}_{t+1},\mathrm{u}_{t+1})=\mathcal{K}\psi(\mathrm{x}_{t},\mathrm{u}_{t})$. Here, a common strategy [32] is to make $\psi(\cdot)$ concatenate the lifted states with the control commands: $\psi(\mathrm{x}_{t},\mathrm{u}_{t})=[\phi(\mathrm{x}_{t});\mathrm{u}_{t}],$ where $\phi(\cdot)$ is the state-dependent lifting function. Finally, we can obtain the new finite approximation as $[\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]=K[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]+r(\mathrm{x}_{t},\mathrm{u}_{t}),$ where $K=\begin{bmatrix}A\in\mathbb{R}^{p\times p}&B\in\mathbb{R}^{p\times m}\\ C\in\mathbb{R}^{m\times p}&D\in\mathbb{R}^{m\times m}\end{bmatrix}$.
Learning Koopman Operator from Data via DMD: The matrix operators $K$ can be inferred from a dataset $\mathcal{D}=[(\mathrm{x}_{1},\mathrm{u}_{1}),\cdots,(\mathrm{x}_{T},\mathrm{u}_{T})]$, which contains the system evolutions and control inputs. Given the choice of observables $\phi(\cdot)$, $K$ is computed by minimizing the residual errors. Specifically, we can obtain $K$ from $\mathcal{D}$ by minimizing the cost function $J(K)$ given below:
$\footnotesize J(K)=\frac{1}{2}\sum_{t=1}^{t=T-1}\|r(\mathrm{x}_{t},\mathrm{u}_{t})\|^{2}=\frac{1}{2}\sum_{t=1}^{t=T-1}\|[\phi(\mathrm{x}_{t+1};\mathrm{u}_{t+1})]-K[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]\|^{2}.$ (2)
Note minimizing $J(K)$ amounts to solving a least-square problem, whose analytical solution can be obtained using the DMD technique [17, 40]: $K=XY^{\dagger}$, where $X=\frac{1}{T-1}\sum_{t=1}^{t=T-1}[\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]\otimes[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]$, and $Y=\frac{1}{T-1}\sum_{t=1}^{t=T-1}[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]\otimes[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]$. Here, $Y^{\dagger}$ denotes the Moore–Penrose inverse of $Y$, and $\otimes$ denotes the outer product. In the standard application of state prediction without predicting the flow of $\mathrm{u}_{t}$, after obtaining the matrix $K$, we can use $\phi(\mathrm{x}_{t+1})=A\phi(\mathrm{x}_{t})+B\mathrm{u}_{t}$ to model the system evolution under control inputs, which facilitates the linear control synthesis [41, 42, 35, 43, 44].
IV Safe Navigation with Learned Koopman Forward Dynamics
This section presents a novel safe navigation framework that leverages the learned Koopman dynamics. A typical model-predictive control (MPC) has been widely adopted as a theoretical backbone for safe navigation. However, the high-level forward dynamics of a legged robot are hard to analyze because it is a closed-loop system with a nonlinear control policy and hybrid contact dynamic, which prevents the existing safety-aware control algorithms from capturing the legged locomotion behaviors accurately.
Instead, we capture a legged robot’s high-level forward dynamics as a linear evolution by adopting the Koopman operator theory, which projects its state space into a high-dimensional feature space. Due to its linearity, we can easily combine the learned dynamics with the existing MPC-based safe navigation frameworks.
IV-A Low-level locomotion controller
Our framework assumes a velocity-tracking controller that takes as input the target velocity and outputs motor commands. Our framework is agnostic to the implementation of a low-level controller. We can use either model-based control theory [14] or a learning-based algorithm, which works best for the given scenario. In our implementation, we train a controller using Proximal Policy Optimization (PPO) inspired by the work of Rudin et al. [1]. We train Unitree’s Aliengo robot using an IsaacGym simulator with 4096 parallel environments for 30 minutes in NVIDIA RTX 4090 GPU until the training converges.
IV-B Data collection for high-level dynamics
We create the dataset of high-level floating base movements by generating 100 locomotion trajectories. Each trajectory consists of 10 seconds of robot state recorded at 50 Hz, which results in 500 data points for each trajectory. We obtain diverse and dynamic trajectories by randomly sampling velocity commands every 0.5 seconds. For each trajectory, we record the high level states of the robot state: $\mathrm{x}=[p_{x},p_{y},\theta,v_{x},v_{y},\omega]$, where $p_{x},p_{y},\theta$ are $x,y$ coordinates and direction heading of the robot and $v_{x},v_{y},\omega$ are $x,y$ linear velocity components and $z$-axis angular velocity of the robot. We also record the velocity tracking commands given to the robot: $\mathrm{u}=[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]$, where the commands are relative to the position and orientation of robot’s state at each step.
For better coverage, we apply a data augmentation technique by applying transformations to the existing dataset. For each episode of length $N$ state transitions $[\mathrm{x}_{0},\mathrm{u}_{0},\dots,\mathrm{u}_{N-1},\mathrm{x}_{N}]$, we extract sequences of length $H$, leading to $(N-H+1)$ sequences per $N$ step trajectory. For each sequence of length $H$, we convert the positional and angular components to the coordinate of its first state.
Formally, for $i$-th state sequence of an episode: $[\mathrm{x}_{i},\dots,\mathrm{x}_{i+H}]$, we convert each state represented in a global coordinate to the local coordinate system with respect to the first state $\mathrm{x}_{i}$. This will lead to a representation of $[\mathrm{x}^{i}_{i},\dots,\mathrm{x}^{i}_{i+H}]$, where $\mathrm{x}^{i}_{j}$ represents a $\mathrm{x}_{j}$ represented in the frame of $\mathrm{x}_{i}$. Similarly for the command sequence $[\mathrm{u}_{i},\dots,\mathrm{u}_{i+H-1}]$, we convert each velocity commands in each step’s robot coordinate to the coordinate of the first step: $[\mathrm{u}^{i}_{i},\dots,\mathrm{u}^{i}_{i+H-1}]$. As a result, we obtain the final dataset $\mathcal{D}$ which contains $2,255,000$ frames of state transitions.
IV-C Learning high-level dynamics
Our main objective in this section is to learn the state transition function $\mathrm{x}_{t+1}=F(\mathrm{x}_{t},\mathrm{u}_{t})$, where $F(\cdot):\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is a high-level dynamics that governs the closed-loop dynamics of nonlinear legged locomotion.
One key challenge in learning $F$ is that it can be arbitrarily nonlinear. One potential way in machine learning to learn the transition function $F$ is to utilize a universal function approximator, a multi-layer perceptron (MLP), to model the input-output mapping of the function $F$ minimizing the $L_{2}$ loss: $||x_{t+1}-\text{MLP}(x_{t},u_{t})||_{2}$ for $\forall t$. However, utilizing MLP dynamics constraint in the MPC framework causes practical challenges, which makes MLP-learned dynamics impractical for general usage. For instance, Direct Multiple Shooting formulation [45, 46, 28], one of the widely used techniques in solving MPC, requires the dynamics constraint containing MLP to match for every time step, making the optimization extremely nonlinear. As a result, nonlinear solvers often fail to solve the optimization with MLP dynamics constraints.
Instead, we use the Koopman operator’s linear evolution to learn the linear evolution in the lifted space, which makes the dynamics constraint in the MPC framework linear. As a result, our approach offers the possibility of adopting linear MPC with control barrier functions. In this work, we use the Koopman forward dynamics as described in Section. III: $\phi(\mathrm{x}_{t+1})=A\phi(\mathrm{x}_{t})+Bu_{t}$, where the state-lifting, $\phi(\cdot):R^{n}\to R^{n^{\prime}}$ $n^{\prime}\geq n$, is a vector-valued lifting function. Inspired by [33], the first $n$ elements of the $\phi(\mathrm{x}_{t})$ are $\mathrm{x}_{t}$ so that we can easily retrieve the original states, which are used for computing obstacle-avoidance constraints. Note that in the case of $n^{\prime}=n$, $\phi(\mathrm{x})$ becomes the same as $\mathrm{x}$, and the solutions $A$ and $B$ becomes linear identification of the high-level dynamics. Finally, we obtain the analytical solution $A$ and $B$ matrices by solving the optimal solution of $J(K)$ in Equation. 2, which in general only takes seconds of cpu-only computation.
IV-D Obstacle avoidance with Koopman MPC
Once we obtain the learned linear evolution in the lifting space, we can incorporate the existing control framework for high-level tasks. In our case, we leverage the existing linear Model-Predictive Control algorithm to design the safety-guaranteed legged navigation controller. Our optimization is formulated as follows:
$\min_{\mathrm{\phi}_{1\cdots H},\mathrm{u}_{1\cdots H}}$ $J(\mathrm{\phi}_{1\cdots H},\mathrm{u}_{1\cdots H})$
s.t. $\phi_{k+1}=A\phi_{k}+B\mathrm{u}_{k}\quad\forall k=0,\dots,H-1$
$h_{i}(\phi_{k})\geq 0\quad\forall h_{i}\in\mathcal{H}\text{ and }\forall k=1,\dots,H$
$\mathrm{u_{k}}\in\mathcal{U}.$
The key idea of our framework is to directly solve lifted variables $\phi_{1\cdots H}$, instead of the original states $\mathrm{x}_{1\cdots H}$, along with control variables $\mathrm{u}_{1\cdots H}$. This formulation is effective because all the dynamics become linear in the lifted space. Therefore, we can easily solve the given optimization using Quadratic Programming.
Our objective function is design to track the given trajectory $\hat{x}_{1\cdots H}$ in the original state spaces:
$J(\mathrm{\phi}_{1\cdots H},\mathrm{u}_{1\cdots H})=\sum_{k=1}^{H}||\mathrm{x_{k}}-\mathrm{\hat{x}_{k}}||_{P}+\sum_{k=0}^{H-1}||\mathrm{u}_{k+1}-\mathrm{u}_{k}||_{2},$
where the $\mathrm{x_{k}}=\mathrm{\phi^{-1}}(\phi_{k})$ for all the timestep $k$. Because the lift function $\phi$ to include the original state information, the inverse operation $\phi^{-1}(\cdot)$ is a linear projection, making the optimization to be quadratic with respect to all the variables.
The first constraint aims to satisfy the learned dynamics. The second constraint is an arbitrary constraint, in our case, an distance barrier function $h_{i}(\phi_{k})=\hat{h}_{i}(\mathrm{x}_{k})=\textbf{dist}(\mathrm{x}_{k},\text{obstacle}_{i})$ for all obstacles. The third constraint confines control signals within the valid region.
Note that this formulation can readily incorporate safety frameworks such as discrete control barrier functions [47], by modifying direct collision checking $\hat{h}(\mathrm{x}_{k})\geq 0$ constraint to that of control barrier function: $\hat{h}(\mathrm{x}_{k+1})\geq\gamma\hat{h}(\mathrm{x}_{k})\quad$ for $\gamma\in[0,1)$, which can still be linear based on obstacle geometry and choice of barrier function $\hat{h}$.
V Experimental Evaluation
We evaluated the our method along with baselines in terms of the long-term forward prediction accuracy (Sec. V-B) and the MPC-based safe navigation performance (Sec. V-C).
V-A Experimental Design
Evaluation Platform: We collected all locomotion data and tested the navigation performance using the Unitree’s Aliengo [48] robot built with IsaacGym simulator [49].
Dynamics Models: We compared various Koopman Dynamics models along with the baselines:
Linear Integrator: An ideal first order linear forward dynamical model: $[p_{x},p_{y},\theta]_{t+1}=[p_{x},p_{y},\theta]_{t}+\frac{dt}{2}[v_{x},v_{y},\omega]_{t}+\frac{dt}{2}[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]_{t}$; $[v_{x},v_{y},\omega]_{t+1}=[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]_{t}$, where $dt=0.02s$ represents the duration of one simulation step. Note that this baseline does not require data for training.
Component-wise Linear Dynamics: Originally proposed in [19], this baseline assumes a linear forward dynamics model for each state component, such as $[p_{x},v_{x}]$ and $\hat{v_{x}}$. Therefore, it has three pairs of matrices $[\bar{A}\in\mathbb{R}^{2\times 2},\bar{B}\in\mathbb{R}^{2\times 1}]$, which are all learned from the dataset $\mathcal{D}$.
Unified Linear Dynamics: This baseline assumes $\phi(\mathrm{x})=\mathrm{x}$ (i.e., no other lifting functions), indicating a linear system evolution under control inputs in the original state space. Similarly, these matrices are learned from the dataset $\mathcal{D}$.
Koopman (P): Inspired by [33], the Koopman model used the lifting function $\phi(\mathrm{x}_{t})$ as polynomials of up to degree three.
Koopman (TD(n)): Another commonly used lifting function is Time-Delay Embedding [50], which concatenates the current system state with the previous $n-1$ history states, i.e., $\phi(\mathrm{x}_{t})=[\mathrm{x}_{t-n+1};\cdots;\mathrm{x}_{t}]$. Note that $\mathrm{u}_{t}$ still remains a single step control command, as required by the MPC formulation. We evaluated different $n$ values: 5, 10, 20, and 30.
V-B Long-term Forward Prediction
Integrator Comp. Linear Unified Linear Koopman (P) Koopman (TD(5)) Koopman (TD(10)) Koopman (TD(20)) Koopman (TD(30))
ValidationSet 0.40($\pm$0.39) 0.27($\pm$0.24) 0.07($\pm$0.09) 0.07($\pm$0.09) 0.07($\pm$0.09) 0.06($\pm$0.09) 0.05($\pm$0.08) 0.04($\pm$0.08)
NewSet1 0.41($\pm$0.49) 0.19($\pm$0.19) 0.09($\pm$0.11) 0.09($\pm$0.11) 0.09($\pm$0.11) 0.09($\pm$0.11) 0.07($\pm$0.10) 0.05($\pm$0.09)
NewSet2 0.39($\pm$0.35) 0.29($\pm$0.26) 0.06($\pm$0.08) 0.06($\pm$0.08) 0.06($\pm$0.07) 0.05($\pm$0.07) 0.04($\pm$0.07) 0.03($\pm$0.06)
NewSet3 0.40($\pm$0.33) 0.32($\pm$0.27) 0.05($\pm$0.07) 0.05($\pm$0.07) 0.05($\pm$0.07) 0.05($\pm$0.07) 0.04($\pm$0.06) 0.03($\pm$0.05)
TABLE I: Average Prediction Error for each dataset, calculated across all 10,000 randomly selected sequences, states, and time steps.
We first evaluate the long-term forward prediction accuracy of each model with respect to the Prediction Error, which measures the absolute differences between predicted and ground-truth states. For a fair comparison, all models were trained on the same dataset, except for the Linear Integrator. We then rolled out the same PPO controller with random velocity commands to collect a new set of 100 locomotion trajectories ($T=500$), which were converted into length-$100$ sequences as the validation dataset.
Figure 3: Average Prediction Error over 10,000 selected sequences. Solid lines represent mean values, and shaded areas indicate standard deviation.
In Fig. 3, we present the average Prediction Error over 10,000 randomly selected sequences for each state at each time step. From the results, it is evident that the Koopman (TD(30)) model significantly outperforms the other models in terms of long-term prediction accuracy. Additionally, its prediction errors are nearly zero within the span of the time-delay embedding. However, surprisingly, the Unified Linear Dynamics and Koopman (P) exhibit very similar performance, as the green and purple lines are nearly indistinguishable, as shown in Fig. 3, suggesting that the polynomial liftings may not effectively capture the underlying dynamics.
We then adjusted the velocity sampling interval (default: 0.5s, used for both training and validation) to 0.1s, 1s, and 3s, and collected three additional test datasets, referred to as NewSet1, NewSet2, and NewSet3. In Table. I, we report each model’s average Prediction Error for each dataset, calculated across all 10,000 randomly selected trajectories, states, and time steps. These results show that the Unified Linear Dynamics and Koopman (P) indeed perform very similarly across all datasets. On the other hand, the results also show that the Koopman forward dynamics benefits from larger time-delay embedding size, and can achieve half the prediction error of the Unified Linear Dynamics when $H=30$. This finding aligns with the results reported in [51], where it shows that providing more history states as input to the neural network policy enhances locomotion performance. However, from the standpoint of practical MPC implementation, there is a trade-off between prediction accuracy and computational burden imposed by larger $A$ and $B$ matrices.
V-C MPC-based Safe Navigation
In this section, we utilize the learned Unified Linear Dynamics model in MPC-based safe navigation settings. We omit Koopman (P) or Koopman (TD) models because Unified Linear Dynamics shows reasonable performance (Fig. 3) and is suitable for real-time control. Note that it is straightforward to extend to the Koopman Dynamics models using either function bases or time-delay embeddings as the lifting function because for both we can easily retrieve $\mathrm{x}_{t}$ from $\phi(\mathrm{x})$ using linear projection for MPC objective and safe constraints.
The navigation framework is tested on three different maps with randomized start and goal selection. To decouple the vision and focus on the effectiveness of our planning framework, we assume the positions and shapes of the obstacles are known. High-level planners are designed to find rough waypoints, which are used as reference trajectories for the MPC controller.
Figure 4: Overview of the navigation framework.
Figure 5: Environments used for Navigation. For corridor 1 and corridor 2, robots are randomly spawned in green start regions and their destination is randomly sampled from blue goal region. For maze environment, start and goal positions are sampled randomly in one of the 5 positions (orange).
Corridor 1 Corridor 2 Maze (width=0.75) Maze (width=0.7)
Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$ Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$ Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$ Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$
Integrator 5.48 1.65 70 5.65 1.25 55 19.68 4.5 90 20.79 12.5 70
Comp. Linear 4.18 1.95 95 4.26 2 100 15.39 4.8 100 14.99 13.4 100
Unified Linear 4.30 1.3 100 4.31 2.25 95 15.21 3.3 100 14.50 6 100
TABLE II: Safe navigation results. Each value represents the mean over 20 random runs for Corridor environments and 10 random runs for Maze environments.
Navigation Environments: We test the effectiveness of our learned models by deploying the robots in 3 different navigation environments: i) corridor 1: a wider corridor with a 70cm gap, ii) corridor 2: a narrower corridor with a 55cm gap, and iii) maze: an artifact maze [28]. For each of the navigation environments, we randomize start and goal positions. We then randomly sample 10 combination of start position and goal position pairs.
High-Level Planner: To navigate complex environments such as mazes, using a search-based high level planner enables creating rough routes toward goals. For each of the environments, we utilize an A* algorithm [52] to create a global path from the start to the goal. At each step of MPC, we truncate the goal and create a local trajectory based on the current position of the robot.
MPC Controller: Based on the desired local trajectory obtained from the high-level planner, we use MPC to compute velocity commands for the low-level controller. We follow the objective in IV-D. Here, we model the geometry of the robot using three circles with a radius of $25$cm, which are equally distributed along the body. The safety constraint is then computed based on the distance between each circle and the polytope modeling of obstacles. The optimization is solved using IPOPT interfaced through CasADi[53]. It’s worth noting that utilizing linear dynamics learned through multi-layer perceptron leads to poor MPC performance due to nonlinear equality constraints.
Experiment results for Linear Integrator, Component-wise Linear, and Unified Linear models are depicted in Table II. We measure i) seconds taken to reach destination, ii) the number of collisions occurred during navigation, and iii) the success rate of the navigation. Here, the success is determined based on whether the robot was able to reach the destination in a desired time window, where we give 6 seconds for corridors and 30 seconds for maze environments. We did not terminate the episode upon minor collisions. However, when the robot fell down and were not be able to recover, we marked them as failure cases.
Overall, the learned Unified Linear model demonstrates the best performance with the lowest number of collisions. This trend becomes more distinctive when it comes to more challenging environments, such as Maze (width=0.75) and Maze (width=0.7), where the Unified Linear model made collisions near 68% or 44% of the second-best method, the Component-wise Linear model. Therefore, we can conclude that the proposed safe navigation framework shows good performance even in complex environments, which have much more dense obstacles compared to the previous works [27].
However, the Linear Integrator model was often unable to succeed the tasks due to significant collisions, making the body fall down. Or, it faced too much deviation from the desired trajectory that it was not able to find adequate commands. Because we did not record additional collisions after the termination, this model achieved the least collisions in Corridor 2 with the worst success rate of $55$%.
It is worth noting that the Component-wise Linear model performed similarly with the Unified Linear model in the corridor tasks. This was a bit surprising because we initially expect much worse navigation performance based on the model prediction performance analyzed in Sec. V-B. Therefore, we suspect that the estimation performance may not be proportionally transferred to the MPC planning performance, depending on the tasks, particularly when they are relatively easy. This leads to room for exploration toward the validity of weak models on safe navigation tasks.
VI Conclusion and Future Work
In this work, we present a novel safe navigation framework for legged robots by combining the Koopman operator theory and the MPC framework. We first effectively learns the dynamics from the collected trajectories using Dynamic Mode Decomposition without the burden of hyperparameter tuning. Then, we develop an MPC framework for safe navigation that efficiently solves the given constrained optimization problem by leveraging the linearity of the learned Koopman dynamics. We show that the Koopman dynamics, especially with Time-delay Embedding, offers improved accuracy for the long-term forward dynamics prediction task. We also deploy our safe navigation framework with the Unified Linear model in environments with narrow passages and mazes and demonstrate improved safety compared to the baselines.
There are several directions for future work. For instance, the utility of the learned linear dynamics is not limited to obstacle avoidance discussed in this work. There exists a wide range of potential tasks, including whole-body control of legged robots, loco-manipulation, or traversing more complex environments. These tasks require modeling the full-body dynamics, which likely requires a careful investigation of the lifting function for the reliable Koopman linearization. Once the Koopman dynamics is learned, it will be possible to leverage similar model-based safety frameworks. We also aim to extend the capability of this method by exploring diverse lifting techniques. One possibility is to utilize neural network-based lifting functions, which could potentially provide a more concise and accurate dynamics model. However, this may introduce complexity to the implementation of the MPC framework. Investigating this trade-off will be an interesting future research direction for legged robots.